OpenStudy (anonymous):

If y varies directly with x and y = -7 when x = -1, find x when y = -84. A. 84 B. 7 C. -12

OpenStudy (anonymous):

it means that y/x = constant

jimthompson5910 (jim_thompson5910):

y = kx -7 = k(-1) -7/(-1) = k 7 = k k = 7 So the equation is y = 7x So when y = -84, then y = 7x -84 = 7x -84/7 = x -12 = x x = - 12